# Constrained Stochastic Recursive Momentum Successive Convex Approximation

### v2

@@ III-C2 Assumption on surrogate functions
Assumptions A6,A7, and A8 restrict the choice of the surrogate functions and have also been used in [19, 14]. Examples of surrogate functions satisfying Assumptions A6-A7 can be found in [45, 19] (see Sec III-D). An implication of Assumptions A3 and A7 is that the function $\phi_{j}({\mathbf{x}}):=g_{j}({\mathbf{x}})-{\tilde{g}}_{j}({\mathbf{x}},{\mathbf{x}}_{t})$ is $2L$-smooth so that
$\phi_{j}({\mathbf{x}})$ $\leq\phi_{j}({\mathbf{x}}_{t})+\langle\nabla\phi_{j}({\mathbf{x}}_{t}),{\mathbf{x}}-{\mathbf{x}}_{t}\rangle+L\left\|{\mathbf{x}}-{\mathbf{x}}_{t}\right\|^{2}=L\left\|{\mathbf{x}}-{\mathbf{x}}_{t}\right\|^{2}$ (17)
@@
Given any $\chi$, it is possible to ascertain $\rho$ by solving the linear programming problem:
@@
In general however, determining global values of the different problem parameters $\sigma$, $L$, $L_{g}$, $\chi$, and $\rho$ may not be easy. In practice, the algorithm parameters are often tuned directly and without first determining the problem parameters.
@@ III-D Examples of Surrogate Functions
The SCA framework allows surrogate functions to be customized for the specific problem at hand. Several examples of surrogates for various types of problems can be found in [19, Sec III-A]. For the sake of completeness, we also mention some of them here.
For $L$-smooth constraint function $g_{j}$, a common choice of surrogate that satisfies all required conditions is the quadratic upper bound ${\tilde{g}}_{j}({\mathbf{x}},{\mathbf{z}})=g_{j}({\mathbf{z}})+\langle\nabla g_{j}({\mathbf{z}}),{\mathbf{x}}-{\mathbf{z}}\rangle+\frac{L}{2}\left\|{\mathbf{x}}-{\mathbf{z}}\right\|^{2}$. If we can write $g_{j}({\mathbf{x}})=g_{j}^{c}({\mathbf{x}})+g_{j}^{\bar{c}}({\mathbf{x}})$ where the first summand is a convex function and the second summand is an $L$-smooth non-convex one, we can use the surrogate ${\tilde{g}}_{j}({\mathbf{x}},{\mathbf{z}})=g_{j}^{c}({\mathbf{x}})+g_{j}({\mathbf{z}})+\langle\nabla g_{j}({\mathbf{z}}),{\mathbf{x}}-{\mathbf{z}}\rangle+\frac{L}{2}\left\|{\mathbf{x}}-{\mathbf{z}}\right\|^{2}$
In addition to these two classes of surrogates, the function $f$ admits a few more. For instance, if $f({\mathbf{x}},{\boldsymbol{\xi}})=f_{1}(f_{2}({\mathbf{x}},{\boldsymbol{\xi}}))$ where $f_{1}$ is a convex function admits the surrogate ${\hat{f}}({\mathbf{x}},{\mathbf{y}},{\boldsymbol{\xi}})=f_{1}(f_{2}({\mathbf{y}},{\boldsymbol{\xi}})+\langle\nabla f_{2}({\mathbf{y}},{\boldsymbol{\xi}}),{\mathbf{x}}-{\mathbf{y}}\rangle)+\frac{\mu}{2}\left\|{\mathbf{x}}-{\mathbf{y}}\right\|^{2}$. In the case of block convex functions, the variable ${\mathbf{x}}$ can be partitioned into disjoint blocks denoted by $\{{\mathbf{x}}_{k}\}_{k=1}^{K}$ and the objective function can be expressed as $f({\mathbf{x}},{\boldsymbol{\xi}})=\sum_{k=1}^{K}f^{i}({\mathbf{x}}_{k},{\mathbf{x}}_{-k},{\boldsymbol{\xi}})$ where each $f^{i}$ is convex in its first argument and ${\mathbf{x}}_{-k}$ collects all variables other than those in the block ${\mathbf{x}}_{k}$. For such functions, we can have the surrogate
${\hat{f}}({\mathbf{x}},{\mathbf{z}},{\boldsymbol{\xi}})=\sum_{i=1}^{I}f^{i}({\mathbf{x}}_{i},{\mathbf{z}}_{-i},{\boldsymbol{\xi}})+\frac{\mu}{2}\left\|{\mathbf{x}}-{\mathbf{z}}\right\|^{2}.$ (20)
@@ IV Complexity analysis
For all $t\geq 1$ and $\beta_{t+1}<\frac{1}{2}$, it holds that
$\frac{1}{16L^{2}}$ ${\mathbb{E}}\left[\frac{\left\|{{\boldsymbol{\varepsilon}}}_{t+1}\right\|^{2}}{\eta_{t}(1-2\beta_{t+1})}\right]\leq\frac{1}{16L^{2}}{\mathbb{E}}\left[\frac{\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}}{\eta_{t}}\right]+\frac{{\mathbb{E}}\left[\eta_{t}\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}\right]}{8}+\frac{1}{8L^{2}}{\mathbb{E}}\left[\frac{\beta_{t+1}^{2}(G_{t+1})^{2}}{\eta_{t}(1-2\beta_{t+1})}\right].$ (27)
@@
If $0<\bar{k}\leq w^{1/3}$, $\tfrac{w^{2/3}}{4\bar{k}^{2}}\geq c\geq 4L^{2}+\frac{G^{2}}{6L\bar{k}^{3}}$, $w\geq\max\{\big{(}\tfrac{2L\bar{k}}{4\mu-3}\big{)}^{3},2G^{2}\}$, and $\mu>\tfrac{3}{4}$, then the following holds
@@
where $M_{T}:=8\tfrac{B_{1}}{{\bar{k}}}+\frac{\sigma^{2}w^{1/3}}{L^{2}\bar{k}^{2}}+\frac{2c^{2}\bar{k}^{2}}{L^{2}}\log(T+2)$ and the ${\tilde{\mathcal{O}}}$ notation hides the dependence on logarithmic terms.
@@
Defining $\Psi_{t}={\mathbb{E}}\left[F({\mathbf{x}}_{t})-F^{\star}\right]+\tfrac{1}{16L^{2}}{\mathbb{E}}[\tfrac{\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}}{\eta_{t-1}(1-2\beta_{t})}]$ and adding (27) with (28), we obtain
$\Psi_{t+1}-\Psi_{t}$ $\leq{\mathbb{E}}\left[\frac{\alpha_{t}}{16L^{2}}\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}\right]-\frac{{\mathbb{E}}\left[\eta_{t}\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}\right]}{8}+\frac{c^{2}}{8L^{2}}{\mathbb{E}}\left[\frac{\eta_{t}^{3}G_{t+1}^{2}}{(1-2c\eta_{t}^{2})}\right],$ (30)
where $\alpha_{t}:=8L^{2}\eta_{t}+\frac{1}{\eta_{t}}-\frac{1}{(1-2\beta_{t})\eta_{t-1}}$. In (30), we observe that the second term is already negative. We also want to select the algorithm parameters such that the first term is also non-positive, leaving the rate depending on the third term, which we will subsequently bound.
@@
$\frac{\bar{k}}{\eta_{t}}-\frac{\bar{k}}{\eta_{t-1}}$ $=(w+\sum_{i=1}^{t}G_{i}^{2})^{1/3}-(w+\sum_{i=1}^{t-1}G_{i}^{2})^{1/3}\leq\frac{G_{t}^{2}}{3(w+\sum_{i=1}^{t-1}G_{i}^{2})^{2/3}}=\frac{G_{t}^{2}\eta_{t-1}^{2}}{3\bar{k}^{2}}$ (31)
Therefore, we can simplify the first two terms in the definition of $\alpha_{t}$ by using (31) and the definition of $\beta_{t}$ as
$\frac{1}{\eta_{t}}-\frac{1}{(1-2\beta_{t})\eta_{t-1}}$ $=\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}-\frac{2\beta_{t}}{(1-2\beta_{t})\eta_{t-1}}$ (32)
$\leq\frac{G_{t}^{2}\eta_{t-1}^{2}}{3\bar{k}^{3}}-\frac{2c\eta_{t-1}}{(1-2\beta_{t})}$ (33)
@@
$\alpha_{t}$ $=\frac{1}{\eta_{t}}-\frac{1}{(1-2\beta_{t})\eta_{t-1}}+8L^{2}\eta_{t}$ (34)
(33)${{\leq}}\frac{G^{2}\eta_{t-1}^{2}}{3\bar{k}^{3}}-2c\eta_{t-1}+8L^{2}\eta_{t-1}$ (35)
$=-2\eta_{t-1}\left(c-4L^{2}-\frac{G^{2}\eta_{t-1}}{6\bar{k}^{3}}\right)\leq-2L^{2}d\eta_{t-1}$ (36)
where $d:=\tfrac{1}{2L^{2}}(c-4L^{2}-\tfrac{G^{2}}{6\bar{k}^{3}})>0$. Therefore the first term on the right of (30) can be dropped.
@@
For the third term on the right of (30), we can use the fact that $w\geq 2G^{2}$ and the concavity of the log function as in [33, Lemma 4] to obtain
@@
$\Psi_{T+1}-\Psi_{1}$ $\leq-\frac{\bar{k}T\Delta_{T}^{2}}{8(w+TG^{2})^{1/3}}+\frac{c^{2}\bar{k}^{3}}{4L^{2}}\ln(T+2)$ (40)
@@
$\Psi_{1}\leq F({\mathbf{x}}_{1})-F^{\star}+\frac{\sigma^{2}}{8L^{2}\eta_{0}}.$ (41)
@@
If $0<\bar{k}\leq w^{1/3}$, $\tfrac{w^{2/3}}{4\bar{k}^{2}}\geq c>4L^{2}+\frac{G^{2}}{6L\bar{k}^{3}}$, and $w\geq\max\{\big{(}\tfrac{2L\bar{k}}{4\mu-3}\big{)}^{3},2G^{2}\}$, and $\mu>\tfrac{3}{4}$, then it also holds that
@@
where $d=\tfrac{1}{2L^{2}}(c-4L^{2}-\tfrac{G^{2}}{6\bar{k}^{3}})$.
The bound in (43a) of Corollary 1 can be established along the lines of Theorem 1 by lower bounding the RHS of (37) and using $\eta_{T}\leq\tfrac{{\bar{k}}}{(w+TG^{2})^{1/3}}$. The bound in (43b) of Corollary 1 can be established on similar lines. Here, instead of dropping the first term on the right of (30), we drop the second term instead, and obtain a bound on the first term by proceeding along similar lines.
Theorem 1 and Corollary 1 establish bounds on the average tracking error and the average progress. These results will later be used to establish the required $\epsilon$-KKT conditions and hence the SFO-complexity bound. To establish these bounds however, we need a strict feasibility condition, which will help us bound the dual variables corresponding to the $t$-th subproblem ($\mathcal{P}_{t}$). This is where the MFCQ condition in Assumption A5 enters into the picture.
@@
The Slater’s condition along with the boundedness assumption in A8 allows us to bound the dual variables for each ($\mathcal{P}_{t}$).
@@
For $t\geq 1$ and ${\tilde{\mathbf{x}}}({\mathbf{x}}_{t})\in{\mathcal{X}}$, if ${\tilde{g}}_{j}({\tilde{\mathbf{x}}}({\mathbf{x}}_{t}),{\mathbf{x}}_{t})\leq-\frac{\rho^{2}}{2L}$ for all $1\leq j\leq J$ and $h_{i}({\tilde{\mathbf{x}}}({\mathbf{x}}_{t}))\leq-\frac{\rho^{2}}{2L}$ for all $1\leq i\leq I$, we have that $\|{\hat{{\boldsymbol{\lambda}}}}_{t}\|_{1}+\|{\hat{{\boldsymbol{\nu}}}}_{t}\|_{1}\leq\frac{2B_{U}L}{\rho^{2}}$.
@@
If $0<\bar{k}\leq w^{1/3}$, $\tfrac{w^{2/3}}{4\bar{k}^{2}}\geq c>4L^{2}+\frac{G^{2}}{6L\bar{k}^{3}}$, $\mu>\tfrac{3}{4}$, and $w\geq\max\{\big{(}\tfrac{2L\bar{k}}{4\mu-3}\big{)}^{3},2G^{2}\}$ with ${\tilde{g}}_{j}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t})\leq-\frac{\rho^{2}}{2L}$ and $h_{i}({{\hat{{\mathbf{x}}}}_{t}})\leq-\frac{\rho^{2}}{2L}$ where $j=1,\ldots,J$ and $i=1,\ldots,I$. Then, there exists some $1\leq t\leq{\tilde{\mathcal{O}}}(\epsilon^{-3/2})$ and ${\hat{{\boldsymbol{\lambda}}}}_{t}\in{\mathbb{R}}^{I}_{+}$, ${\hat{{\boldsymbol{\nu}}}}_{t}\in{\mathbb{R}}^{J}_{+}$ such that the point $({{\hat{{\mathbf{x}}}}_{t}},{\hat{{\boldsymbol{\lambda}}}}_{t},{\hat{{\boldsymbol{\nu}}}}_{t})$ is $\epsilon$-KKT optimal.
@@
$\nabla$ ${\tilde{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})$
$=\nabla{\hat{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})+(1-\beta_{t})({\mathbf{z}}_{t}-\nabla f({\mathbf{x}}_{t-1},{\boldsymbol{\xi}}_{t}))$ (47)
@@
Likewise, since ${\hat{f}}(\cdot,{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})$ is $L$-smooth from Assumption A6, it follows that
$\left\|\nabla{\hat{f}}({\mathbf{x}}_{t},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})-\nabla{\hat{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})\right\|\leq L\left\|{\boldsymbol{\delta}}_{t}\right\|$. Also from (18) and since ${\hat{{\boldsymbol{\lambda}}}}_{t}\geq 0$, we have that
@@
$\left\|\Big{(}\nabla{\mathbf{g}}({{\hat{{\mathbf{x}}}}_{t}})-\nabla{\tilde{{\mathbf{g}}}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t})\Big{)}{\hat{{\boldsymbol{\lambda}}}}_{t}\right\|\leq\sum_{j=1}^{J}\left\|[{\hat{{\boldsymbol{\lambda}}}}_{t}]_{j}(\nabla g_{j}({{\hat{{\mathbf{x}}}}_{t}})-\nabla{\tilde{g}}_{j}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t}))\right\|\leq 2L\left\|{\hat{{\boldsymbol{\lambda}}}}_{t}\right\|_{1}\left\|{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t}\right\|.$ (53)
@@
$\left\|{\boldsymbol{\pi}}_{t}\right\|$ $\leq 2L(1+\tfrac{2B_{U}L}{\rho^{2}})\left\|{\boldsymbol{\delta}}_{t}\right\|+\left\|{\boldsymbol{\varepsilon}}_{t}\right\|$ (54)
@@
$\leq\left(\left(2L+\frac{4B_{U}L^{2}}{\rho^{2}}\right)^{2}+\frac{1}{d}\right)\frac{2M_{T}(w+TG^{2})^{1/3}}{T}$
$={\tilde{\mathcal{O}}}(T^{-2/3})$ (56)

### v3

@@ III-C2 Assumption on surrogate functions
Assumptions A8,A9, and A10 restrict the choice of the surrogate functions and have also been used in [21, 15]. Examples of surrogate functions satisfying Assumptions A8-A9 can be found in [47, 21] (see Sec III-D). An implication of Assumptions A3 and A9 is that the function $\phi_{j}({\mathbf{x}}):=g_{j}({\mathbf{x}})-{\tilde{g}}_{j}({\mathbf{x}},{\mathbf{x}}_{t})$ is $2L$-smooth so that
$\phi_{j}({\mathbf{x}})$ $\leq\phi_{j}({\mathbf{x}}_{t})+\langle\nabla\phi_{j}({\mathbf{x}}_{t}),{\mathbf{x}}-{\mathbf{x}}_{t}\rangle+L\left\|{\mathbf{x}}-{\mathbf{x}}_{t}\right\|^{2}$
$=L\left\|{\mathbf{x}}-{\mathbf{x}}_{t}\right\|^{2}$ (21)
@@
Given any $\chi$, it is possible to ascertain $\rho$ by solving the optimization problem:
@@
$\left\|{\mathbf{d}}_{{\tilde{\mathbf{x}}}}\right\|_{2}\leq 1$
In general however, determining global values of the different problem parameters $\sigma$, $L$, $\chi$, and $\rho$ may not be easy. In practice, the algorithm parameters are often tuned directly and without first determining the problem parameters.
@@ III-D Examples of Surrogate Functions
The SCA framework allows surrogate functions to be customized for the specific problem at hand. Several examples of surrogates for various types of problems can be found in [21, Sec III-A]. For the sake of completeness, we also mention some of them here.
For $L$-smooth constraint function $g_{j}$, a common choice of surrogate that satisfies all required conditions is the quadratic upper bound ${\tilde{g}}_{j}({\mathbf{x}},{\mathbf{z}})=g_{j}({\mathbf{z}})+\langle\nabla g_{j}({\mathbf{z}}),{\mathbf{x}}-{\mathbf{z}}\rangle+\frac{L}{2}\left\|{\mathbf{x}}-{\mathbf{z}}\right\|^{2}$. If we can write $g_{j}({\mathbf{x}})=g_{j}^{c}({\mathbf{x}})+g_{j}^{\bar{c}}({\mathbf{x}})$ where the first summand is a convex function and the second summand is an $L$-smooth non-convex one, we can use the surrogate ${\tilde{g}}_{j}({\mathbf{x}},{\mathbf{z}})=g_{j}^{c}({\mathbf{x}})+g_{j}({\mathbf{z}})+\langle\nabla g_{j}({\mathbf{z}}),{\mathbf{x}}-{\mathbf{z}}\rangle+\frac{L}{2}\left\|{\mathbf{x}}-{\mathbf{z}}\right\|^{2}$.
In addition to these two classes of surrogates, the function $f$ admits a few more structures. For instance, if $f({\mathbf{x}},{\boldsymbol{\xi}})=f_{1}(f_{2}({\mathbf{x}},{\boldsymbol{\xi}}))$ where $f_{1}$ is a convex function admits the surrogate ${\hat{f}}({\mathbf{x}},{\mathbf{y}},{\boldsymbol{\xi}})=f_{1}(f_{2}({\mathbf{y}},{\boldsymbol{\xi}})+\langle\nabla f_{2}({\mathbf{y}},{\boldsymbol{\xi}}),{\mathbf{x}}-{\mathbf{y}}\rangle)+\frac{\mu}{2}\left\|{\mathbf{x}}-{\mathbf{y}}\right\|^{2}$. In the case of block convex functions, the variable ${\mathbf{x}}$ can be partitioned into disjoint blocks denoted by $\{{\mathbf{x}}_{k}\}_{k=1}^{K}$ and the objective function can be expressed as $f({\mathbf{x}},{\boldsymbol{\xi}})=\sum_{k=1}^{K}f^{k}({\mathbf{x}}_{k},{\mathbf{x}}_{-k},{\boldsymbol{\xi}})$ where each $f^{k}$ is convex in its first argument and ${\mathbf{x}}_{-k}$ collects all variables other than those in the block ${\mathbf{x}}_{k}$. For such functions, we can have the surrogate
${\hat{f}}({\mathbf{x}},{\mathbf{z}},{\boldsymbol{\xi}})=\sum_{k=1}^{K}f^{k}({\mathbf{x}}_{k},{\mathbf{z}}_{-k},{\boldsymbol{\xi}})+\frac{\mu}{2}\left\|{\mathbf{x}}-{\mathbf{z}}\right\|^{2}.$ (24)
@@ IV Complexity analysis
For all $t\geq 1$ and $\beta_{t+1}<\frac{1}{4}$, it holds that
$\frac{1}{16L^{2}}$ ${\mathbb{E}}\left[\frac{\left\|{{\boldsymbol{\varepsilon}}}_{t+1}\right\|^{2}}{\eta_{t}}\right]\leq\frac{1}{16L^{2}}{\mathbb{E}}\left[\frac{(1-\beta_{t+1})\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}}{\eta_{t}}\right]+\frac{{\mathbb{E}}\left[\eta_{t}\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}\right]}{8}+\frac{1}{8L^{2}}{\mathbb{E}}\left[\frac{\beta_{t+1}^{2}(G_{t+1})^{2}}{\eta_{t}}\right].$ (31)
@@
Under Assumptions A1-A5 and A8, if $0<\bar{k}\leq w^{1/3}$, $\tfrac{w^{2/3}}{4\bar{k}^{2}}\geq c\geq 8L^{2}+\frac{G^{2}}{3\bar{k}^{3}}$, $w\geq\max\{\big{(}\tfrac{2L\bar{k}}{4\mu-3}\big{)}^{3},2G^{2}\}$, and $\mu>\tfrac{3}{4}$, then the following holds
@@
where $M_{T}:=8\frac{B_{1}}{{\bar{k}}}+\frac{\sigma^{2}w^{1/3}}{2L^{2}\bar{k}^{2}}+\frac{c^{2}\bar{k}^{2}}{L^{2}}\log(T+2)$, and the ${\tilde{\mathcal{O}}}$ notation hides the dependence on logarithmic terms.
@@
Defining $\Psi_{t}={\mathbb{E}}\left[F({\mathbf{x}}_{t})-F^{\star}\right]+\tfrac{1}{16L^{2}}{\mathbb{E}}[\tfrac{\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}}{\eta_{t-1}}]$ and adding (31) with (32), we obtain
$\Psi_{t+1}-\Psi_{t}$ $\leq{\mathbb{E}}\left[\frac{\alpha_{t}}{16L^{2}}\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}\right]-\frac{{\mathbb{E}}\left[\eta_{t}\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}\right]}{8}+\frac{c^{2}}{8L^{2}}{\mathbb{E}}\left[\eta_{t}^{3}G_{t+1}^{2}\right],$ (34)
where $\alpha_{t}:=8L^{2}\eta_{t}+\frac{(1-\beta_{t+1})}{\eta_{t}}-\frac{1}{\eta_{t-1}}$. In (105), we observe that the second term is already negative. We also want to select the algorithm parameters such that the first term is also non-positive, leaving the rate depending on the third term, which we will subsequently bound.
@@
$\frac{\bar{k}}{\eta_{t}}-\frac{\bar{k}}{\eta_{t-1}}$ $=(w+\sum_{i=1}^{t}G_{i}^{2})^{1/3}-(w+\sum_{i=1}^{t-1}G_{i}^{2})^{1/3}$
$\leq\frac{G_{t}^{2}}{3(w+\sum_{i=1}^{t-1}G_{i}^{2})^{2/3}}=\frac{G_{t}^{2}\eta_{t-1}^{2}}{3\bar{k}^{2}}$ (35)
Therefore, we can simplify the last two terms in the definition of $\alpha_{t}$ by using (35) and the definition of $\beta_{t}$ as
$\frac{(1-\beta_{t+1})}{\eta_{t}}-\frac{1}{\eta_{t-1}}$ $=\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}-\frac{\beta_{t+1}}{\eta_{t}}$ (36)
$\leq\frac{G_{t}^{2}\eta_{t-1}^{2}}{3\bar{k}^{3}}-c\eta_{t}$ (37)
@@
$\alpha_{t}$ $=\frac{(1-\beta_{t+1})}{\eta_{t}}-\frac{1}{\eta_{t-1}}+8L^{2}\eta_{t}$ (38)
(37)${{\leq}}\frac{G^{2}\eta_{t-1}^{2}}{3\bar{k}^{3}}-c\eta_{t-1}+8L^{2}\eta_{t-1}$ (39)
$=-2\eta_{t-1}\left(\frac{c}{2}-4L^{2}-\frac{G^{2}\eta_{t-1}}{6\bar{k}^{3}}\right)\leq-2L^{2}d\eta_{t-1}$ (40)
where $d:=\tfrac{1}{2L^{2}}(c-8L^{2}-\tfrac{G^{2}}{3\bar{k}^{3}})>0$. Therefore the first term on the right of (105) can be dropped.
@@
For the third term on the right of (105), we can use the fact that $w\geq 2G^{2}$ and the concavity of the log function as in [20, Lemma 4] to obtain
@@
$\Psi_{T+1}-\Psi_{1}$ $\leq-\frac{\bar{k}T\Delta_{T}^{2}}{8(w+TG^{2})^{1/3}}+\frac{c^{2}\bar{k}^{3}}{8L^{2}}\ln(T+2)$ (44)
@@
$\Psi_{1}\leq F({\mathbf{x}}_{1})-F^{\star}+\frac{\sigma^{2}}{16L^{2}\eta_{0}}.$ (45)
@@
Under Assumptions A1-A5 and A8, along with the conditions of Theorem 1, it also holds that
@@
where $M_{T}:=8\frac{B_{1}}{{\bar{k}}}+\frac{\sigma^{2}w^{1/3}}{2L^{2}\bar{k}^{2}}+\frac{c^{2}\bar{k}^{2}}{L^{2}}\log(T+2)$, and $d=\tfrac{1}{2L^{2}}(c-8L^{2}-\tfrac{G^{2}}{3\bar{k}^{3}})$.
The bound in (47a) of Corollary 1 can be established along the lines of Theorem 1 by lower bounding the RHS of (41) and using $\eta_{T}~{}\geq~{}\tfrac{{\bar{k}}}{(w+TG^{2})^{1/3}}$. The bound in (47b) of Corollary 1 can be established on similar lines. Here, instead of dropping the first term on the right of (105), we drop the second term instead, and obtain a bound on the first term by proceeding along similar lines.
It is important to note that in Lemma 1, the last term of (31) affects the upper bounds, causing the convergence rate to depend on a logarithmic term. However, if the adaptive step sizes in (11) are replaced with fixed step sizes $\eta=\tfrac{{\bar{k}}}{T^{1/3}}$ and $\beta=c\eta^{2}$, the upper bound in Theorem 1 improves to $\mathcal{O}(T^{-2/3})$. Building on this observation, we present the following corollary, whose proof is omitted for brevity.
Corollary 2.
Under Assumptions A1-A2, A4-A6 and A8, if $0<\bar{k}\leq 1$, $\tfrac{1}{4\bar{k}^{2}}\geq c>8L^{2}$, and $\mu>\tfrac{3}{4}$, with $\eta=\frac{{\bar{k}}}{T^{1/3}}$ and $\beta=c\eta^{2}$, it holds that
$\Delta_{T}\leq\sqrt{\frac{M^{\prime}_{T}}{T^{2/3}}}={\mathcal{O}}\left(\frac{1}{T^{1/3}}\right)$ (48a)
$\tfrac{1}{T}\sum_{t=1}^{T}{\mathbb{E}}\left[\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}\right]\leq\frac{M^{\prime}_{T}}{T^{2/3}}={\mathcal{O}}\left(\frac{1}{T^{2/3}}\right)$ (48b)
$\tfrac{1}{T}\sum_{t=1}^{T}{\mathbb{E}}\left[\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|\right]\leq\sqrt{\frac{\bar{M^{\prime}}_{T}}{d^{\prime}T^{2/3}}}={\mathcal{O}}\left(\frac{1}{T^{1/3}}\right)$ (48c)
where $M^{\prime}_{T}:=8\frac{B_{1}}{{\bar{k}}}+\frac{\sigma^{2}}{2L^{2}\bar{k}^{2}}+\frac{c^{2}\bar{k}^{2}}{L^{2}}G^{2}$, and $d^{\prime}=\tfrac{1}{2L^{2}}(c-8L^{2})$
The proof of Corollary 2 follows on the lines of Theorem 1 and has been detailed in Appendix E.
Theorem 1 and Corollary 1 establish bounds on the average tracking error and the average progress. These results will later be used to establish the required $\epsilon$-KKT conditions and hence the SFO-complexity bound. To establish these bounds however, we need a strict feasibility condition, which will help us bound the dual variables corresponding to the $t$-th sub-problem ($\mathcal{P}_{t}$). This is where the MFCQ condition in Assumption A7 enters into the picture.
@@
The Slater’s condition along with the boundedness assumption in A10 allows us to bound the dual variables for each ($\mathcal{P}_{t}$).
@@
Under conditions of Lemma 4, for $t\geq 1$, we have that $\|{\hat{{\boldsymbol{\lambda}}}}_{t}\|_{1}+\|{\hat{{\boldsymbol{\nu}}}}_{t}\|_{1}\leq\frac{2B_{U}L}{\rho^{2}}$.
@@
Under Assumptions A1-A5 and A7-A10, along with the conditions of Theorem 1, Corollary 1 and Lemma 4, there exists some $1\leq t\leq{\tilde{\mathcal{O}}}(\epsilon^{-3/2})$ and ${\hat{{\boldsymbol{\lambda}}}}_{t}\in{\mathbb{R}}^{I}_{+}$, ${\hat{{\boldsymbol{\nu}}}}_{t}\in{\mathbb{R}}^{J}_{+}$ such that the point $({{\hat{{\mathbf{x}}}}_{t}},{\hat{{\boldsymbol{\lambda}}}}_{t},{\hat{{\boldsymbol{\nu}}}}_{t})$ is $\epsilon$-KKT optimal.
@@
$\nabla{\tilde{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})$ $=\nabla{\hat{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})+(1-\beta_{t})({\mathbf{z}}_{t}-\nabla f({\mathbf{x}}_{t-1},{\boldsymbol{\xi}}_{t}))$ (52)
@@
Likewise, since ${\hat{f}}(\cdot,{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})$ is $L$-smooth from Assumption A8, it follows that $\left\|\nabla{\hat{f}}({\mathbf{x}}_{t},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})-\nabla{\hat{f}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})\right\|\leq L\left\|{\boldsymbol{\delta}}_{t}\right\|$. Also from (22) and since ${\hat{{\boldsymbol{\lambda}}}}_{t}\geq 0$, we have that
@@
$\left\|\Big{(}\nabla{\mathbf{g}}({{\hat{{\mathbf{x}}}}_{t}})-\nabla{\tilde{{\mathbf{g}}}}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t})\Big{)}{\hat{{\boldsymbol{\lambda}}}}_{t}\right\|$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\leq\sum_{j=1}^{J}\left\|[{\hat{{\boldsymbol{\lambda}}}}_{t}]_{j}(\nabla g_{j}({{\hat{{\mathbf{x}}}}_{t}})-\nabla{\tilde{g}}_{j}({{\hat{{\mathbf{x}}}}_{t}},{\mathbf{x}}_{t}))\right\|\leq 2L\left\|{\hat{{\boldsymbol{\lambda}}}}_{t}\right\|_{1}\left\|{{\hat{{\mathbf{x}}}}_{t}}-{\mathbf{x}}_{t}\right\|.$ (58)
@@
$\left\|{\boldsymbol{\pi}}_{t}\right\|$ $\leq 2L(1+\tfrac{2B_{U}L}{\rho^{2}})\left\|{\boldsymbol{\delta}}_{t}\right\|+\left\|{\boldsymbol{\varepsilon}}_{t}\right\|.$ (59)
@@
$\leq\left(M_{T}\left(2L+\frac{4B_{U}L^{2}}{\rho^{2}}\right)^{2}+\frac{\bar{M}_{T}}{d}\right)\frac{2(w+TG^{2})^{1/3}}{T}={\tilde{\mathcal{O}}}(T^{-2/3})$ (61)